. A group of students was asked to indicate their preference for regular or diet soft drinks. of 14 males, 11 preferred regular soft drinks. Of 11 females, 7 preferred diet soft drinks. Find the probability that a student selected at random is a: (a) diet soft drinks or is a female: (b) regular soft drinks or is a female: (c) regular soft drinks or is a male
Answer #1
Male Female Total
Regular soft drink 11 4 15
Diet soft drink 3 7 10
Total 14 11 25

(a)

P(Diet soft drink OR Female) = P(Diet soft drink) + P(Female) - P(Diet soft drink & Female)

                                           = 10/25 + 11/25 - 7/25 = 14/25 = 0.56

So,

Answer is

0.56

(b)

P(Regular soft drink OR Female) = P(Regular soft drink) + P(Female) - P(Regular soft drink & Female)

                                                 = 15/25 + 11/25 - 4/25 = 22/25 = 0.88

So,

Answer is:

0.88

(c)

P(Regular soft drink OR Male) = P(Regular soft drink) + P(Male) - P(Regular soft drink & Male)

                                               = 15/25   + 14/25 - 11/25 = 18/25 = 0.72

So,

Answer is:

0.72